this is an Integer's question...
40-(30-3)
33-(3-13)
8-9+2
8(-9)+2
2(-1)-3(4)

Answer by checkley77(12844)

40-(30-3)
40-27=13 ans.
33-(3-13)
33-(-10)=33+10=43 ans.
8-9+2
10-9=1 ans.
8(-9)+2
-72+2=-70 ans.
2(-1)-3(4)
-2-12=-14 ans,